OpenStudy (anonymous):

For the function defined by f(x)=3x^2+5x+2, find the following values. (a) f(-4)= (b) f(-x)= (c) -f(x)= (d) f(x+h)=

OpenStudy (mathstudent55):

For (a), (b), and (d), all you need to do is to replace x in the polynomial with each of the values you are given, and evaluate. (a) f(-4) = 3*(-4)^2 + 5(-4) + 2 (b) f(-x) = 3*(-x)^2 + 5(-x) + 2 (d) f(x + h) = 3*(x + h)^2 + 5(x + h) + 2 For (b), just take the negative of the polynomial. -f(x) = -(3x^2 + 5x + 2)